# BST: Badminton Stroke-type Transformer for Skeleton-based Action Recognition in Racket Sports

### v2

@@ Abstract
Badminton, known for having the fastest ball speeds among all sports, presents significant challenges to the field of computer vision, including player identification, court line detection, shuttlecock trajectory tracking, and player stroke-type classification. In this paper, we introduce a novel video segmentation strategy to extract frames of each player’s racket swing in a badminton broadcast match. These segmented frames are then processed by two existing models: one for Human Pose Estimation to obtain player skeletal joints, and the other for shuttlecock trajectory detection to extract shuttlecock trajectories. Leveraging these joints, trajectories, and player positions as inputs, we propose Badminton Stroke-type Transformer (BST) to classify player stroke-types in singles. To the best of our knowledge, experimental results demonstrate that our method outperforms the previous state-of-the-art on the largest publicly available badminton video dataset, ShuttleSet, which shows that effectively leveraging ball trajectory is likely to be a trend for racket sports action recognition. The code is publicly available at https://github.com/Va6lue/BST-Badminton-Stroke-type-Transformer.
@@ 1 Introduction
Fortunately, in badminton singles, the shuttlecock trajectory can serve as an interactive medium between the two players. To the best of our knowledge, TemPose [16] is the first model to incorporate shuttlecock trajectory information as an auxiliary input. However, we argue that the shuttlecock trajectory plays an even more crucial role in recognizing badminton stroke actions. Consider a hypothetical scenario where the shuttlecock is completely removed from a badminton match video, leaving only the two players performing air swings. From a human perspective, would it still be possible to accurately determine the type of stroke performed based solely on the players’ movements? This highlights the indispensable role of shuttlecock trajectory information in a robust badminton stroke recognition model. Therefore, we treat the shuttlecock trajectory as a primary input in our model. The main contributions of this paper are summarized as follows:
  • We introduce a novel badminton video segmentation strategy to extract frames that have high relevance to the target badminton stroke swung by the player.
  • We propose Badminton Stroke-type Transformer (BST) models that outperform the previous state-of-the-art on the largest publicly available badminton video dataset, ShuttleSet [32].
  • We show that effectively leveraging shuttlecock trajectory information can significantly improve the performance of badminton stroke-type classification.
@@ 3 Method
In this section, we present the key components of our method. In Sec. 3.1, we present a video segmentation strategy which segments a match video into clips. Each clip contains two or three strokes. In Sec. 3.2, we introduce a data preprocessing work which produces player poses, positions and shuttlecock trajectories from stroke clips. In Sec. 3.3, we present a Transformer-based architecture which takes poses as inputs and produces classification results.
3.1 Video Segmentation Strategy
@@
for $j=1,2,\ldots,n^{(i)}$ and $1\leq i\leq r$. In Eq. 1, from the perspective of having comprehensive player pose information, $t$ ensures that the target hit frame $F_{h^{(i)}_{j}}$ is at the middle of the stroke clip $C^{(i)}_{j}$ and is ideally set to half of the interval between every two consecutive hit frame numbers. However, it is a fixed parameter, and clearly, all stroke clips defined in Eq. 1 realistically contain at least one hit frame and possibly many hit frames.
Intuitively, consecutive strokes are highly correlated, and correlation between far-away strokes drops very rapidly. Thus, to increase the stroke classification accuracy, we would like to provide additional strokes to the classifier. On the other hand, we would not like to provide too many strokes to the classifier, as far-away strokes are much less correlated with, and can be even irrelevant to, the target stroke being classified. Our goal is for each stroke clip to include information only from the previous and next strokes swung by the opponent. Achieving this goal is difficult with a fixed-width clipping strategy using a fixed $t$. For this reason, we propose a variable-width clipping strategy as follows. First, define
$\hat{h}^{(i)}_{j-1}$ =$=\begin{cases}h^{(i)}_{j-1},&\text{if }j>1\\ h^{(i)}_{j}-t,&\text{if }j=1,\end{cases}$
$\hat{h}^{(i)}_{j+1}$ =$=\begin{cases}h^{(i)}_{j+1}+\epsilon,&\text{if }j<n^{(i)}\\ h^{(i)}_{j}+t,&\text{if }j=n^{(i)},\end{cases}$ (2)
@@
If the target stroke being classified is the first or last stroke in a rally, we set the lower or upper bound of the stroke clip to be the same value as used in Eq. 1. In Eq. 2, $\epsilon$ is a small parameter less than $t$. Looking at this strategy from the perspective of shuttlecock trajectory, it captures clean and complete trajectories in three stages: (1) the shuttlecock flying towards the target player’s court, (2) the shuttlecock trajectory after being hit by the target player, and (3) the direction of the shuttlecock trajectory after the opponent’s response. These stages respectively represent (1) the target player’s observation and reaction phase, (2) the actual trajectory of the target stroke, and (3) the opponent’s response pattern. This allows the model to not only judge based on (2) but also understand the target player’s common strategies after going through (1) and infer the target stroke-type back from (3). Thus, our goal is achieved, and the most informative segment of the shuttlecock trajectory related to the target stroke is captured. The reason that the stroke information in (3) is less than that in (1), which means that $\epsilon$ cannot be set too large, is to preserve the characteristic of keeping the target hit frame close to the middle of the stroke clip under the ideal condition in Eq. 1.
[…]
In practice, we also set some limits on $\hat{h}^{(i)}_{j-1}$ and $\hat{h}^{(i)}_{j+1}$ to prevent the model from accessing irrelevant information caused by remaining unclean data. Specifically, the actual time distance of these two frame numbers from the target frame number does not exceed 1.5 seconds (before adding $\epsilon$). Additionally, $t$ is set to the half of the FPS of the video representing 0.5 seconds interval, and $\epsilon$ is set to $\frac{t}{2}$.
@@ 3.2 Model Inputs Extracted from Clips
For the player poses and positions, we extract human poses from the clips using MMPose [8] toolbox at first. The models we utilized in MMPose are RTMPose [18] for 2D pose estimation and MotionBERT [38] for further 3D pose estimation. (We choose 2D poses for the input of the model. See Sec. 5 for more details.) However, there are more than two people, including the two players, spectators, and referees, in the clips. To extract the exact two player poses, we need the court lines to determine the player positions to filter out the irrelevant poses. Fortunately, the court lines information is available in our dataset (Sec. 4.1). Even if the court lines are not available, we can use the algorithm proposed in MonoTrack [23] to extract the court lines.
@@ 3.3 Badminton Stroke-type Transformer (BST)
The architecture of BST is derived from a Transformer-based model TemPose [16], which, to our knowledge, is the state-of-the-art badminton stroke classification model. A key distinction is that BST shifts its focus from player poses to shuttlecock trajectory information. Below, we describe several versions of BST, as follows: BST-0 serves as a prototype for the BST series. BST-0 takes player poses and shuttlecock trajectory information as its inputs, whereas BST-1, BST-2, and BST-3 incorporate additional player position information.
Figure 1: Architecture of BST-0
@@ 3.3.1 BST-0
Initially, the player poses222 In practice: Top player pose is Pose 1; Bottom player pose is Pose 2. and shuttlecock positions are processed through a TCN module. After the outputs from this TCN are passed through a Transformer Encoder, individual class tokens with spatiotemporal representations are obtained, as shown in Fig. 1.
The Transformer modules in this paper are more similar to that used in TemPose [16] than the original Transformer [31]. At the beginning, a learnable class token is prepended to the input sequence for summarization, and a learnable positional embedding is added to the input as well. We can formulate our Transformer Encoder as follows:
@@
where $X^{(l)}\in\mathbb{R}^{S\times D_{model}}$ is the output to the $l$-th layer, $S$ is the sequence length, $D_{model}$ is the dimension used in the model, LN denotes the layer normalization, MHSA is the multi-head self-attention mechanism, and FFN is the feedforward network. Before passing through the MHSA, the input $X^{(l-1)}$ which simplifies here to $X$ is split into:
@@
where $W_{O}\in\mathbb{R}^{(D_{A}\times H)\times D_{model}}$ represents the linear transformation from the dimenstion $D_{A}$ to the original model dimension $D_{model}$.
@@
Next, a key concept of our approach is ensuring that the model focuses on extracting critical information from the shuttlecock trajectory. To achieve this, we design a Cross Transformer layer with a Multi-Head Cross Attention (MHCA), which differs from MHSA in one key aspect: in the attention mechanism, $K$ and $V$ are derived from the shuttlecock trajectory latent, while $Q$ is generated from the player pose latent or the player position latent, respectively, to perform cross-attention. As a result, the subsequent outputs follow two distinct pathways. One pathway is dedicated to the trajectory information related to the first player poses, while the other pathway focuses on the trajectory information related to the second player poses. Both pathways are fed into a second Transformer Encoder. Finally, the model generates logits through an MLP (multi-layer perceptron) head, which are converted into probability values using the softmax function.
@@
Trans1 and Trans2 are inspired by TemPose [16] (Temporal and Interactional Transformer), but Trans2 is actually different from their Interactional Transformer that serves people as a sequence. Our Trans2 still processes the frame sequence to obtain the class token containing the trajectory information related to player 1 or 2. For CrossTrans, our idea involves calculating the similarity between pose and trajectory, and using these similarity values to pick the value arrays in $V$ (from trajectory), which is a critical information from our perspective, an insight gained after observing our experimental results in Sec. 4 (TemPose-PF vs. TemPose-SF in Tab. 1).
In practice, there are 2 layers in Trans1, 1 layer in CrossTrans, and 1 layer in Trans2.
3.3.2 Pose Postion Fusion (PPF)
In subsequent models, player position information was incorporated as an additional input to the model. $X$ is preprocessed by a Pose Position Fusion (PPF) module to generate a fused representation of pose and position, as shown in Fig. 2. The output of the MLP in the PPF module multiplies back with the original poses and becomes the coefficients on the player poses.
Figure 2: Pose Postion Fusion (PPF)
3.3.3 BST-1
@@
In BST-1 shown in Fig. 3, a Clean Gate module is designed to refine the information derived from the shuttlecock trajectory by leveraging the interaction strength between the two players and the shuttlecock. This helps the model filter out noise in the trajectory caused by the opponent’s strokes. The Clean Gate module shown in Fig. 4 takes three inputs: (1) the class token representing shuttlecock information, (2) the fusion class token of the shuttlecock with the player one, and (3) the fusion class token of the shuttlecock with the player two. The input (2) and (3) are processed through min pooling and a MLP to generate values, which are then used to adjust the shuttlecock information class token, input (1), accordingly.
Figure 3: Architecture of BST-1
Figure 4: Clean Gate
3.3.4 BST-2
@@
In BST-2, shown in Fig. 5, we compare the cosine similarity between input (1) and (2) with that between input (1) and (3). The difference between these two values gives $\tilde{\alpha}$. Since cosine similarity ranges from -1 to 1, $\tilde{\alpha}\in[-2,2]$. We normalize it using $\alpha=(\tilde{\alpha}+2)/4$ so that $\alpha\in[0,1]$. $\alpha$ denotes the coefficient on the player one’s information, so $1-\alpha$ is the coefficient on the player two’s information. This ensures that the player with a stronger correlation to the overall shuttlecock trajectory exerts a greater influence on the input of the MLP head.
Figure 5: Architecture of BST-2
3.3.5 BST-3
BST-3, as shown in Fig. 6, combines the benefits of BST-1, which focuses on purifying shuttlecock information, and BST-2, which emphasizes the information derived from the target player.
Figure 6: Architecture of BST-3

### v3

@@ Abstract
Badminton, known for having the fastest ball speeds among all sports, presents significant challenges to the field of computer vision, including player identification, court line detection, shuttlecock trajectory tracking, and player stroke-type classification. In this paper, we introduce a novel video clipping strategy to extract frames of each player’s racket swing in a badminton broadcast match. These clipped frames are then processed by three existing models: one for Human Pose Estimation to obtain human skeletal joints, another for shuttlecock trajectory tracking, and the other for court line detection to determine player positions on the court. Leveraging these data as inputs, we propose Badminton Stroke-type Transformer (BST) to classify player stroke-types in singles. To the best of our knowledge, experimental results demonstrate that our method outperforms the previous state-of-the-art on the largest publicly available badminton video dataset (ShuttleSet), another badminton dataset (BadmintonDB), and a tennis dataset (TenniSet). These results suggest that effectively leveraging ball trajectory is a promising direction for action recognition in racket sports.
@@ 1 Introduction
Fortunately, in badminton singles, the shuttlecock trajectory can serve as an interactive medium between the two players. To the best of our knowledge, TemPose [18] is the first model to incorporate shuttlecock trajectory information as an auxiliary input. However, we argue that the shuttlecock trajectory plays an even more crucial role in recognizing badminton stroke actions. Consider a hypothetical scenario where the shuttlecock is completely removed from a badminton match video, leaving only the two players performing air swings. From a human perspective, would it still be possible to accurately determine the type of stroke performed based solely on the players’ movements? Moreover, players may deliberately perform deceptive movements to confuse their opponents, but the shuttlecock trajectory always reflects the actual stroke-type and contains no misleading information. These points highlight the indispensable role of shuttlecock trajectory information in a robust badminton stroke recognition model. Therefore, we treat the shuttlecock trajectory as a primary input in our model. The main contributions of this paper are summarized as follows:
  • We introduce a novel badminton video clipping strategy to extract frames that have high relevance to the target badminton stroke swung by the player.
  • We propose Badminton Stroke-type Transformer (BST) architecture, including its backbone BST-0 and several enhancing modules, that outperforms the previous state-of-the-art on the largest publicly available badminton video dataset (ShuttleSet [34]), another badminton dataset (BadmintonDB [3]), and a tennis dataset (TenniSet [12]).
  • We show that effectively leveraging ball trajectory information (when designing the method) can significantly improve the performance of badminton stroke classification, and even tennis stroke classification.
  • Our BST series models achieve faster training speed compared to previous state-of-the-art methods, as detailed in the supplementary material.
  • We present a detailed analysis and comparison of model performance using 2D and 3D joints in the supplementary material.
@@ 3 Method
In this section, we present the key components of our method. In Sec. 3.1, we present a video clipping strategy which segments a match video into clips. In Sec. 3.2, we introduce a data preprocessing work which produces player poses, positions and shuttlecock trajectories from stroke clips. In Sec. 3.3, we present a Transformer-based architecture which takes those as inputs and produces classification results.
3.1 Video Clipping Strategy
(a) Fixed-width strategy
(b) Complete-pose strategy
(c) Our strategy (from the perspective of pose)
(d) Our strategy (from the perspective of trajectory)
Figure 1: The idea of video clipping strategies. The black player is the target player, and the red player is his/her opponent.
@@ 3.1 Video Clipping Strategy
for $j=1,2,\ldots,n^{(i)}$ and $1\leq i\leq r$, as shown in LABEL:sub@fig:clipping_strategy-a. The parameter, $t$, in this strategy ensures that the target hit frame $F_{h^{(i)}_{j}}$ is in the middle of the stroke clip $C^{(i)}_{j}$. Ideally, $t$ is set to half the interval between two consecutive hit frames. However, since $t$ is a fixed value, this presents a dilemma: a small value for $t$ may result in a clip that is too short, failing to capture complete player pose information, while a large value may include extraneous information (too many other hit frames) that is too far from the target stroke. To address this issue, we need a different approach. We can solve this by using the previous and next hit frame numbers from the opponent, as shown in LABEL:sub@fig:clipping_strategy-b. Therefore, a complete-pose clip is obtained.
We further extend this approach. Intuitively, consecutive strokes are highly correlated. Therefore, our goal is for each stroke clip to contain information pertaining not only to the target stroke, but also to the opponent’s previous and next strokes. To achieve this, we propose our strategy as shown in LABEL:sub@fig:clipping_strategy-c and LABEL:sub@fig:clipping_strategy-d. We just need to add a small value $\epsilon$, which is less than $t$, to the next hit frame number $h^{(i)}_{j+1}$ from the opponent. As we mentioned in Sec. 1, the shuttlecock trajectory is an indispensable piece of information for stroke-type classification. This small value, therefore, plays a crucial role in capturing partial trajectory for the opponent’s next stroke, which is not included in the complete-pose strategy. Comprehensively, these kinds of stroke clips can be thought of as three stages:
  1. 1.
    Containing the second half of the pose and complete trajectory from the opponent’s previous stroke, and the first half of the pose for the target stroke.
  2. 2.
    Containing the second half of the pose and complete trajectory for the target stroke, and the first half of the pose from the opponent’s next stroke.
  3. 3.
    Containing the pose beginning in the second half of the opponent’s next stroke, including its partial trajectory.
Thus, our goal is achieved. This allows the downstream model to not only judge based on the actual target trajectory in stage 2 and understand the target player’s common strategies after going through stage 1, but also infer the target stroke-type back from stage 3. The reason that stage 3 is shorter than stage 1, which means that $\epsilon$ cannot be set too large, is to preserve the characteristic of keeping the target hit frame close to the middle of the stroke clip under the ideal condition in the fixed-width strategy. In other words, the complete target pose information should be placed in the central area of the stroke clip (LABEL:sub@fig:clipping_strategy-c). Furthermore, the target trajectory information in stage 2 is also now nearly centrally located, as demonstrated in LABEL:sub@fig:clipping_strategy-d. In formal terms, we can define
$\hat{h}^{(i)}_{j-1}$ $=\begin{cases}h^{(i)}_{j-1},&\text{if }j>1\\ h^{(i)}_{j}-t,&\text{if }j=1,\end{cases}$
$\hat{h}^{(i)}_{j+1}$ $=\begin{cases}h^{(i)}_{j+1}+\epsilon,&\text{if }j<n^{(i)}\\ h^{(i)}_{j}+t,&\text{if }j=n^{(i)},\end{cases}$ (3)
@@
If the target stroke being classified is the first or last stroke in a rally, we set the lower or upper bound of the stroke clip to be the same value as used in Eq. 2.
@@ 3.2 Model Inputs Extracted from Clips
For the player poses and positions, we extract human poses from the clips using MMPose [9] toolbox at first. The models we utilized in MMPose are RTMPose [20] for 2D pose estimation and MotionBERT [41] for further 3D pose estimation. (We choose 2D poses for better downstream accuracy.) However, there are more than two people, including the two players, spectators, and referees, in the clips. To extract the exact two player poses, we need the court lines to determine the player positions to filter out the irrelevant poses. The court lines information is available in one of our datasets. For others not available, we used the algorithm proposed in MonoTrack [25] or a deep learning model TennisCourtDetector [36] to extract the court lines.
@@ 3.3 Badminton Stroke-type Transformer (BST)
Figure 2: Architecture of BST. (In practice: Top player is blue, and Bottom player is green.)
The architecture of BST is derived from a Transformer-based model TemPose [18], which, to our knowledge, is the state-of-the-art badminton stroke classification model. A key distinction is that BST shifts its focus from player poses to shuttlecock trajectory information. The following subsections and Fig. 2 describe the components of BST in detail.
@@ 3.3.1 BST-0
BST-0 serves as the backbone of the BST architecture. Initially, the player poses and shuttlecock positions are each processed through dedicated TCN modules. After the outputs from this TCN are passed through a Transformer Encoder with two Transformer Layers, individual class tokens representing spatiotemporal features are obtained.
The Transformer Encoder modules in this paper are more similar to that used in TemPose [18] than the original Transformer [33]. At the beginning, a learnable class token is prepended to the input sequence for conclusion, and a learnable positional embedding is added to the input as well. We can formulate a Transformer Layer as follows:
@@
where $X^{(l)}\in\mathbb{R}^{S\times D_{model}}$ is the output to the $l$-th layer in this Transformer Encoder, $S$ is the sequence length, $D_{model}$ is the dimension used in the model, LN denotes the layer normalization, MHSA is the multi-head self-attention mechanism, and FFN is the feedforward network. Before passing through the MHSA, the input $X^{(l-1)}$ which simplifies here to $X$ is split into:
@@
where $W_{O}\in\mathbb{R}^{(D_{A}\times H)\times D_{model}}$ represents the linear transformation from the attention dimension $D_{A}$ to the original model dimension $D_{model}$.
@@
Besides the different module used on handling the poses at the beginning from TemPose [18], a key concept of our approach is ensuring that the model focuses on extracting critical information from the shuttlecock trajectory. To achieve this, we design a Cross Transformer Layer (CrossTrans) with a Multi-Head Cross Attention (MHCA), which differs from MHSA in one key aspect: in the attention mechanism, $K$ and $V$ are derived from the shuttlecock trajectory latent, while $Q$ is generated from the player pose latent or the player position latent, respectively, to perform cross-attention. As a result, the subsequent outputs are the shuttlecock trajectory latents belonging to the two players, respectively. Then, both latents are fed into a second Transformer Encoder. After that, the model leverages the overall shuttlecock information, the latent related to the blue player, and the latent related to the green player to generate logits through an MLP (multi-layer perceptron) Head. Finally, a softmax function is applied to the outputs of the MLP Head to obtain probability values.
@@
Trans1 and Trans2 are inspired by TemPose [18] (Temporal and Interactional Transformer), but Trans2 is actually different from their Interactional Transformer that serves people as a sequence. Our Trans2 still processes the frame sequence to obtain the class token containing the trajectory information related to these two players, respectively. For CrossTrans, our idea involves calculating the similarity between pose and trajectory, and using these similarity values to pick the value arrays in $V$ (from trajectory), which is a critical information from our perspective, an insight gained after observing our experimental results in Sec. 4 (TemPose-PF vs. TemPose-SF in Tab. 1).
3.3.2 Pose Position Fusion (PPF) Module
This module generates a fused representation of pose and position. An MLP within the module outputs coefficients, which are multiplied with the original poses. This process imbues each joint with the influence of its corresponding court position. The model, BST, is composed of BST-0 and this module.
3.3.3 Clean Gate (CG) Module
@@
This module is designed to refine the information derived from the shuttlecock trajectory by leveraging the interaction strength between the two players and the shuttlecock. This helps the model filter out noise in the trajectory caused by the opponent’s strokes. The model, BST-CG, is composed of BST and this module, and the overall trajectory information before passing through MLP Head is replaced by the cleaned trajectory information after passing through this module.
Model Publication Modality SP $P$ Acc Macro-F1 Min-F1 Acc-2
ST-GCN [35] AAAI 2018 J × × 0.7758 0.7352 0.3726 0.9200
BlockGCN [40] CVPR 2024 J × × 0.7652 0.7298 0.3788 0.9164
SkateFormer [10] ECCV 2024 J × × 0.7710 0.7363 0.4355 0.8900
ProtoGCN [24] CVPR 2025 J × × 0.7746 0.7306 0.3326 0.9176
TemPose-V [18] CVPRW 2023 J × × 0.7730 0.7360 0.4052 0.9248
TemPose-V [18] CVPRW 2023 J+B × × 0.7756 0.7408 0.4286 0.9266
TemPose-PF J+B × 0.7942 0.7704 0.4912 0.9384
TemPose-SF J+B × 0.8100 0.7808 0.4872 0.9418
BST-0 J+B × 0.8194 0.7924 0.5210 0.9444
TemPose-TF [18] CVPRW 2023 J+B 0.8189 0.7943 0.4928 0.9496
BST J+B 0.8206 0.7952 0.5331 0.9499
BST-CG J+B 0.8210 0.7954 0.5296 0.9481
BST-AP J+B 0.8229 0.7992 0.5532 0.9484
BST-CG-AP J+B 0.8254 0.7983 0.5196 0.9503
Table 1: Model comparison on ShuttleSet (25 classes) with fixed-width strategy. J and B in the modality denote the players’ joints and bones. SP and $P$ represent additional inputs of shuttlecock and player positions. TemPose-PF and TemPose-SF are intermediate products derived by us from the original model, positioned between TemPose-V and TemPose-TF. The suffix ”F” denotes the fusion of the additional inputs. Acc-2 denotes top-2 accuracy.
Model Modality SP $P$ Acc Macro-F1 Min-F1 Acc-2
TemPose-SF J+B × 0.8171 0.7936 0.5186 0.9515
BST-0 J+B × 0.8284 0.8041 0.5822 0.9584
TemPose-TF [18] J+B 0.8235 0.8028 0.5423 0.9567
BST J+B 0.8287 0.8049 0.5523 0.9602
BST-CG J+B 0.8307 0.8048 0.5444 0.9597
BST-AP J+B 0.8288 0.8056 0.5652 0.9586
BST-CG-AP J+B 0.8322 0.8097 0.5762 0.9594
Table 2: Model comparison on ShuttleSet (25 classes) with our clipping strategy.
3.3.4 Aim Player (AP) Module
@@
We compare the cosine similarity between the overall trajectory information and each player’s trajectory information. The difference between these two values yields $\tilde{\alpha}$. Since cosine similarity ranges from -1 to 1, $\tilde{\alpha}\in[-2,2]$. We normalize it using $\alpha=(\tilde{\alpha}+2)/4$ so that $\alpha\in[0,1]$. $\alpha$ denotes the coefficient on the blue player’s information, so $1-\alpha$ is the coefficient on the green player’s information. This ensures that the player with a stronger correlation to the overall shuttlecock trajectory exerts a greater influence on the input to the MLP Head. The model, BST-AP, is composed of BST and this module but without passing the overall trajectory information through MLP Head. This is because we have a final model, BST-CG-AP, which combines both CG and AP modules with BST.